OpenStudy (anonymous):

–x + 3y = 9 into slope intercept form? is it -x + -9/3 = y?

OpenStudy (austinl):

Slope intercept form is y = mx + b

OpenStudy (anonymous):

@austinL Yes, I know. but with working the problem out and stuff is it the way i wrote it up there?

OpenStudy (austinl):

-x + 3y = 9 3y = x + 9 y = 1/3x + 3 \[y=\frac{1}{3}x +3\]